OpenStudy (anonymous):

Walk me through it How many grams of zinc metal will react completely with 5.4 liters of 2.8 M HCl? Show all of the work needed to solve this problem. Zn (s) + 2 HCl (aq) ---> ZnCl2 (aq) + H2 (g)

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4f6be83ee4b0772daa077864 i hope it helped...

OpenStudy (somy):

do u get the working?

OpenStudy (anonymous):

No I wanted to wait for you haha you make it make sense

OpenStudy (chmvijay):

would you like to walk with me through this ?

OpenStudy (anonymous):

Sure

OpenStudy (chmvijay):

ok from equation can u tell me how many mole of Zn reacts with how many moles of HCl?

OpenStudy (anonymous):

2.8 moles of HCI right?

OpenStudy (chmvijay):

where does 2.8 come from? i didnt find it from the equation

OpenStudy (somy):

oh

OpenStudy (anonymous):

"How many grams of zinc metal will react completely with 5.4 liters of 2.8 M HCl?"

OpenStudy (chmvijay):

ok @Somy came better she know how well she can make u unserstnd s this :P

OpenStudy (anonymous):

Okay thank you!

OpenStudy (chmvijay):

bbye

OpenStudy (somy):

lol where are we?

OpenStudy (anonymous):

Well all I know is that there are 2.8 moles of HCI

OpenStudy (somy):

how did u get it?

OpenStudy (anonymous):

I literally says it in the equation

OpenStudy (anonymous):

"How many grams of zinc metal will react completely with 5.4 liters of 2.8 M HCl?"

OpenStudy (somy):

How many grams of zinc metal will react completely with 5.4 liters of 2.8 M HCl? Show all of the work needed to solve this problem. Zn (s) + 2 HCl (aq) ---> ZnCl2 (aq) + H2 (g)

OpenStudy (somy):

2.8 M M is molarity, thus its CONCENTRATION not mole

OpenStudy (anonymous):

Ohhhh

OpenStudy (somy):

yeah so for HCl you are given molarity = 2.8 M and volume = 5.4 L

OpenStudy (somy):

Molarity and concentration are same things thus units are same volume we always convert it to dm3 right? you are given in Liters Liters is equal to dm3 so these two are also same thing

OpenStudy (somy):

get it?

OpenStudy (anonymous):

Okay so if it's already in Liters I don't have to touch it

OpenStudy (somy):

yes exactly :)

OpenStudy (somy):

what is the formula of molarity?

OpenStudy (anonymous):

M = moles solute/liters solution

OpenStudy (somy):

yup so you need to find mole right?

OpenStudy (anonymous):

Yeah

OpenStudy (somy):

good so find it now

OpenStudy (anonymous):

2.8/5.4 = 0.51

OpenStudy (somy):

wrong

OpenStudy (somy):

hmmm u still have problems with making subjects in formulas

OpenStudy (somy):

i think we should have a small tutorial on this later

OpenStudy (anonymous):

Okay

OpenStudy (somy):

conc= mole/dm3 if you make mole as subject that will be mole= conc*dm3

OpenStudy (anonymous):

Okay 2.8*5.4 = 15.12

OpenStudy (somy):

yup that's right

OpenStudy (somy):

Zn (s) + 2 HCl (aq) ---> ZnCl2 (aq) + H2 (g) okay so mole of 2HCl is 15.12moles what is mole of Zn?

OpenStudy (somy):

look at the ratio between HCl and Zn in reaction its 1:2 right?

OpenStudy (anonymous):

I hate ratios but I see that

OpenStudy (somy):

yeah so if 2 -> 15.12 mole 1 -> X

OpenStudy (anonymous):

7.56!

OpenStudy (somy):

bingo :)

OpenStudy (somy):

so Zn= 7.56 mole

OpenStudy (somy):

now tell me what is mole formula?

OpenStudy (anonymous):

n = m/M

OpenStudy (somy):

yup and u need to find mass right?

OpenStudy (somy):

u know Mr of Zn from periodic table and u already calculated mole

OpenStudy (anonymous):

Whats the formula for that?

OpenStudy (somy):

molecular mass

OpenStudy (somy):

formula for what?

OpenStudy (anonymous):

wait all I have to do is look of the molecular mass of Zn right

OpenStudy (anonymous):

which is 65.

OpenStudy (anonymous):

n = m/M n = 65/ 7. 56 = 8. 59

OpenStudy (somy):

noo

OpenStudy (somy):

mole= mass/ Mr

OpenStudy (somy):

so mass = mole*Mr

OpenStudy (anonymous):

7.56*65 = 491. 4

OpenStudy (somy):

yup

OpenStudy (anonymous):

Now we do mole = mole/Mr?

OpenStudy (somy):

no lol we are done :)

OpenStudy (somy):

you got the mass of zinc they asked you for :)

OpenStudy (anonymous):

Thank you:)

OpenStudy (somy):

you are welcome :) i'll open a new question where i'll teach you how to work with formulas ok?

OpenStudy (anonymous):

Okay!

OpenStudy (somy):

cool :)